OpenStudy (anonymous):

7f+g=5, for f just answer

OpenStudy (anonymous):

All you have to do is subtract g on both sides and divide by coefficient of f. Let me know what you get!

OpenStudy (anonymous):

f=5/7+g/7 ?

Directrix (directrix):

7f+g=5 7f = 5 - g (1/7) (7f) = (1/7) (5 - g) f = (5 - g)/7 = 5/7 - g/7

OpenStudy (anonymous):

No! It's f=5-g/7 You need to subtract g and divide by 7.

OpenStudy (anonymous):

you need to divide 5 by 7 too

OpenStudy (anonymous):

Yup but if you need to safe a bit of lead, you can put the whole thing on top of 7 (: